OpenStudy (anonymous):

y= x^2+4x-5 y=2x-2 please help.

OpenStudy (anonymous):

equate both

OpenStudy (anonymous):

Can you please show me some examples?

OpenStudy (anonymous):

if y =a and y = b then a = b

OpenStudy (anonymous):

I'll give a try...thank you.

OpenStudy (anonymous):

u will need to solve a quadratic equation

OpenStudy (anonymous):

Yes, So far I have a, no solution.

OpenStudy (anonymous):

nah it wil hv a solution

OpenStudy (anonymous):

show me what u did

OpenStudy (anonymous):

Well... I graphed it on a graphing calculator.

OpenStudy (anonymous):

what equation did u graph?

OpenStudy (anonymous):

y=x^2+4x-5 y=2x-2

OpenStudy (anonymous):

x2 + 4x - 5 = 2x - 2 x2 +2x -3 = 0 x2 +3x - x - 3 = 0 (x + 3)(x - 1) = 0 So, x = 1 Or x = -3 is the solution..

OpenStudy (anonymous):

that means x^2+4x-5 =2x-2 isnt it?

OpenStudy (anonymous):

Yes, the same you said...

OpenStudy (anonymous):

yes. Thank you for your help.